OpenStudy (anonymous):

Is the reciprocal of 2/sin b = 2cscb?

OpenStudy (anonymous):

Well it is true that\[ \frac{1}{\sin (x)} = \csc(x) \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Does that answer your question?

OpenStudy (anonymous):

ok. well I guess I didn't realize how it broke down into a reciprocal, like the actual breakdown of it>?

OpenStudy (anonymous):

Start with\[ \frac{1}{\sin (x)} = \csc(x) \]Multiply both sides by \(2\) \[ \frac{2}{\sin (x)} = 2\csc(x) \]Let \(x=b\) \[ \frac{2}{\sin (b)} = 2\csc(b) \]

OpenStudy (anonymous):

ahhhh!!!thank you. greatly appreciated.

Parth (parthkohli):

Pardon me, but the reciprocal of \(\dfrac{2}{\sin(b)} = \dfrac{\sin(b)}{2}\). :)